OpenStudy (anonymous):

Kobe's overtime pay is $5 an hour more than his regular pay. He worked 8 hours at his regular wage and 3 hours at his overtime wage. He earned $114. What is Kobe's regular wage per hour?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

8x+3(y+5) = 114 8x+3y+15 = 114 8x+3y = 99 72+27 = 99 Regular wage = 9

OpenStudy (anonymous):

overtime=5+r regular=r 3(5+r)+8r=114 15+3r+8r=114 15+11r=114 11r=114-15 11r=99 r=99/11 r=9